OpenStudy (anonymous):

Which of the following statements best describes the effect of replacing the graph of f(x) with the graph of f(x) - 3? The graph shifts 3 units up. The graph shifts 3 units down. The graph shifts 3 units left. The graph shifts 3 units right.

OpenStudy (anonymous):

you can always this on an equation. for example f(x)=x^2  and f(x)-3=x^2 is 

OpenStudy (anonymous):

Shifts up?

OpenStudy (anonymous):

yes

OpenStudy (lena772):

A